OpenStudy (anonymous):

Whats the absolute value of 11/4 (thats 11 over 4) and how did you find the answer?

jimthompson5910 (jim_thompson5910):

The absolute value of any positive number is that exact same positive number So \[\large \left|\frac{11}{4}\right|=\frac{11}{4}\]

jimthompson5910 (jim_thompson5910):

Remember: the absolute value of a number is the distance that number is from zero

OpenStudy (anonymous):

so if 11/4 was negative it would still be 11/4?

jimthompson5910 (jim_thompson5910):

yes \[\large \left|-\frac{11}{4}\right|=\frac{11}{4}\]

OpenStudy (anonymous):

Kinda like the true distance from zero either way